TRUTH TABLES: MAKE A TRUTH TABLE FOR EACH OF THE FOLLOWING STATEMENT


1. p ∧ q 3 points



2. p ∨ q 3 points



3. p∨~q 4 points



1
Expert's answer
2021-01-29T09:17:37-0500

Let us make a truth table for each of the following statement:


1. "p \land q"


"\ \ \begin{array}{||c|c||c||}
\hline\hline
p & q & p\land q\\
\hline\hline
F & F & F\\
\hline
F & T & F\\
\hline
T & F & F\\
\hline
T & T & T\\
\hline\hline
\end{array}"


2. "p \lor q"


"\ \ \begin{array}{||c|c||c||}
\hline\hline
p & q & p\lor q\\
\hline\hline
F & F & F\\
\hline
F & T & T\\
\hline
T & F & T\\
\hline
T & T & T\\
\hline\hline
\end{array}"



3. "p\lor\sim q"


"\ \ \begin{array}{||c|c||c|c||}
\hline\hline
p & q & \sim q& p\lor \sim q\\
\hline\hline
F & F & T & T\\
\hline
F & T & F & F\\
\hline
T & F & T & T\\
\hline
T & T & F & T\\
\hline\hline
\end{array}"